OpenStudy (anonymous):

the polynomial 0.041h-0.018a-2.69 can be used to estimate the lung capacity, in liters, of a female with height h, in centimeters, and age a in years. Find the lung capacity of a 20 year old woman who is 170 centimeters tall?

OpenStudy (jlastino):

h= 170cm a=20 0.041(170)-0.018(20)-2.69 = lung capacity 6.97 - 0.36 = 6.61 6.61 -2.69 =3.92 3.92 liters

OpenStudy (anonymous):

What Jlastino did was simply filling in the values in the polynomial you gave. There are 2 variables: h and a. You have the h: 170cm and you have the a: 20 years. Now substitute those values in your polynomial and perform the calculation.